# Discrete Cosine Transform Based Decorrelated Attention for Vision Transformers

### v3

@@ IV-C ImageNet-1K Classification
IV-D Ablation Study
IV-D1 Multiple DCT Initialization
Table IV presents the ablation study about initializing multiple matrices of $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ as the DCT matrix. Although such an initialization still improves the accuracy for Swin-T and Swin-S, the improvement is smaller than with the single DCT initialization. For example, in the Swin-T experiments, if $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$ are both initialized as the DCT matrix, the top-1 accuracy is 81.540%, lower than 81.568% from initializing $\mathbf{W}_{Q}$ only and 81.728% from initializing $\mathbf{W}_{K}$ only. This deterioration is caused by the weight symmetry problem [6] from initializing the multiple weights as the same values. In other words, the attention mechanism is designed to attend to relevant information in the input sequence. If $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are the same, the model may not effectively capture the relationships between different positions in the input sequence. Moreover, one of the advantages of self-attention is its ability to model complex dependencies between different positions in the sequence. If $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are the same, the capacity of the model to capture diverse patterns will be limited, and the performance of the model will drop correspondingly.
TABLE IV: ImageNet-1K ablation study on multiple weights initialization as the DCT matrix.
Model DCT Top-1 Acc. Top-5 Acc.
Swin-T - 81.474% 95.776%
Swin-T-DCT Q 81.568% 95.862%
Swin-T-DCT K 81.728% 95.862%
Swin-T-DCT V 81.528% 95.850%
Swin-T-DCT QK 81.540% 95.754%
Swin-T-DCT QV 81.548% 95.748%
Swin-T-DCT KV 81.536% 95.770%
Swin-S - 83.196% 96.360%
Swin-S-DCT Q 83.298% 96.298%
Swin-S-DCT K 83.378% 96.398%
Swin-S-DCT QK 83.208% 96.378%
IV-D2 Non-Trainable DCT Weights Initialization
In this case, computing queries, keys, and values is equivalent to shifting the DCT representation of the input patterns. Table V shows we obtain comparable accuracy results as the baseline but use fewer trainable parameters on the CIFAR-10 dataset. It also shows on the ImageNet-1K dataset if we take non-trainable DCT weights to compute $\mathbf{Q}$ or $\mathbf{K}$, the top-1 accuracy drops insignificantly (less than 0.3%). It is worthwhile noting that in this work, the DCT is implemented by matrix multiplication. Hence, the computational overhead of using non-trainable DCT weights is the same as the vanilla attention. Still, the DCT with butterfly operation can further reduce this cost as we only need to apply DCT in each attention layer.
TABLE V: Ablation study on non-trainable DCT weights.
@@ IV-D2 Non-Trainable DCT Weights Initialization
IV-D3 Removing DCT in DCT-Compressed Attention
This section investigates the necessity of DCT in the DCT-compressed attention. As Table VI presents, when we remove the DCT in the DCT-compressed attention and retain other components, the accuracy of the revised Swin-T models on ImageNet-1K decreases across all compression ratios. These experiments show that DCT is essential because it transforms feature representations into the frequency domain, reducing redundancy and enhancing the model’s ability to capture essential information. As a result, the attention mechanism benefits from improved expressiveness and robustness, even under compression. The accuracy drop observed in Table VI confirms that DCT plays a crucial role in preserving discriminative features, demonstrating its necessity in the proposed approach.
TABLE VI: ImageNet-1K ablation study on removing DCT in DCT-compressed attention.
Model Params (M) Top-1 Acc. Top-5 Acc.
Swin-T 28.29 81.474% 95.776%
Swin-T-DCT-0.25 22.21 79.580% 94.936%
Swin-T-DCT-0.25 (Remove DCT) 22.21 78.710% 94.638%
Swin-T-DCT-0.5 23.43 80.762% 95.520%
Swin-T-DCT-0.5 (Remove DCT) 23.43 80.276% 95.086%
Swin-T-DCT-0.75 25.45 81.474% 95.728%
Swin-T-DCT-0.75 (Remove DCT) 25.45 81.064% 95.468%
IV-E COCO 2017 Object Detection and Instance Segmentation
We further evaluate the proposed methods on COCO 2017 object detection and instance segmentation tasks [55]. We use the Swin Transformers with ImageNet-1K pre-trained weights as the backbone and the Cascade Mask R-CNN [65, 66] as the pipeline. Since initializing $\mathbf{W}_{K}$ as the DCT matrix obtains the best accuracy results in the ImageNet-1K experiments, we use Swin-T-DCT-K and Swin-S-DCT-K to evaluate the DCT-based initialization. As Table VII presents, both Swin-T-DCT-K and Swin-S-DCT-K obtain slightly better average precision (AP) results than the vanilla models in the two tasks. Swin-T-DCT-0.75 obtains comparable AP results as Swin-T with about 3M fewer parameters, and Swin-S-DCT-0.75 gets better AP results than Swin-S with about 5M fewer parameters.
TABLE VII: COCO results. Our methods with higher or lower $\text{AP}^{\text{Mask}}$ are highlighted in pink or cyan.
Backbone Params (M) Object Detection Instance Segmentation
$\text{AP}^{\text{Mask}}$ $\text{AP}^{\text{Mask}}_{0.5}$ $\text{AP}^{\text{Mask}}_{0.75}$ $\text{AP}^{\text{Mask}}$ $\text{AP}^{\text{Mask}}_{0.5}$ $\text{AP}^{\text{Mask}}_{0.75}$
Swin-T [17] 86 50.3 69.1 54.3 43.7 66.6 47.3
DCFormer-T [36] 86 43.5 62.6 47.4 38.0 59.3 40.7
@@ IV-E COCO 2017 Object Detection and Instance Segmentation
V Conclusion
In this work, we introduced two new methods using Discrete Cosine Transform (DCT) to decorrelate the input in the attention for vision transformers. First, we presented a novel DCT-based initialization approach to address the challenges associated with training attention weights. We demonstrated a significant improvement over traditional initialization, providing a robust foundation for the attention mechanism. Second, we formulated a DCT-based compression method that can effectively reduce the computational overhead by truncating high-frequency components. It resulted in smaller weight matrices that compromised considerable accuracy. We validated the efficacy of these breakthroughs, showcasing their potential to enhance the performance and efficiency of Transformer models in the classification tasks. These contributions underscore the importance of thoughtful initialization strategies and compression techniques in advancing the capabilities of the Swin Transformer architectures.
VI Acknowledgment
This work was supported by National Science Foundation (NSF) grants 1934915, 2229659, U.S. Department of Energy (DOE) grant DE-SC0023715, and National Institutes of Health (NIH) grant R01-CA240639.

### v5

@@ 4.2 Image Classification Benchmark
Table 3: ImageNet-1K validation dataset evaluation.
ImageNet-1K
Table 3 presents the performance comparison on the ImageNet-1K dataset. Same as on the CIFAR-10 dataset, DCT-based initialization consistently improves performance on both Swin-T and Swin-S models, regardless of which attention weight was initialized as the DCT matrix. For example, when we initialized the weights to compute keys ($\mathbf{K}$) as the DCT weights, the accuracy of Swin-T was improved from 81.474% to 81.728%, and Swin-S was increased from 83.196% to 83.378%, respectively. We then evaluated the DCT-compressed attention with different truncating rates and compared the results with DCFormer Li et al. (2023), a DCT-based approach designed to reduce computational cost. When 75% of DCT coefficients were kept, we obtained a comparable accuracy result on Swin-T and a slightly better accuracy result on Swin-S, while reducing the trainable parameters from 28.29M to 25.45M (13.1% off) and from 49.61M to 44.45M (10.4% off), respectively. Flops were reduced from 4.49G to 4.18G (6.9% off) for Swin-T and 8.13G to 8.74G (7.0% off) for Swin-S. Retaining fewer coefficients yields further savings at the cost of accuracy. In contrast, DCFormer-SW-T/S with $\tau=2$ ($4\times$ frequency compression) substantially reduces FLOPs but suffers notable performance degradation, as it treats low- and high-frequency components equally. We further validated our methods on the ViT-B-32 backbone Dosovitskiy et al. (2020). DCT-based initialization of the key projection improves accuracy from 73.120% to 73.614%. Applying DCT-compressed attention with 75% coefficient retention further improves accuracy to 73.226%, while reducing parameters from 88.22M to 75.83M (14.0%) and FLOPs from 4.41G to 4.05G (8.2%).
(a) Initializing multiple attention weights with the DCT matrix.
(b) Removing DCT from compressed attention.
Figure 4: Ablation study on ImageNet-1K for (a) initializing multiple attention weights with the DCT matrix and (b) removing DCT from DCT-compressed attention. Swin-T serves as the backbone model. In (a), “Q”, “K”, and “V” indicate which projection weights are initialized with the DCT matrix. Combinations such as “QK” denote that multiple weights (e.g., query and key) are initialized using the DCT matrix.
4.3 Ablation Study
Non-Trainable DCT Weights Initialization
This study investigates the effectiveness of DCT weights as a principled starting point for attention projection initialization. Specifically, we fix the projection weights as DCT matrices, under which the computation of queries, keys, and values becomes equivalent to applying fixed shifts in the DCT representation of the input features. Table 4 shows that our method achieved accuracy comparable to the baseline on the CIFAR-10 dataset while requiring fewer trainable parameters. On the ImageNet-1K dataset, non-trainable DCT weights for computing $\mathbf{Q}$ or $\mathbf{K}$ led to only a marginal drop in top-1 accuracy (less than 0.3%). In this work, the DCT was implemented via matrix multiplication, incurring the same computational cost as standard attention; however, this overhead can be further reduced using butterfly-based DCT implementations.
@@ 4.3 Ablation Study
Table 4: Ablation study on non-trainable DCT weights.
Multiple DCT Initializations
This study shows why we only initialize one instead of multiple projection weights. Fig. 4a presents an ablation study on initializing multiple projection matrices with the DCT matrix. Although this approach still improved top-1 accuracy for Swin-T, the gains were reduced compared to initializing only a single matrix. For example, initializing both $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$ resulted in a top-1 accuracy of 81.540%, lower than 81.568% by initializing only $\mathbf{W}_{Q}$ and 81.728% by initializing only $\mathbf{W}_{K}$. This degradation was likely due to the weight symmetry problem Hu et al. (2019a), which often arises when multiple weight matrices are initialized with identical values. In the context of attention mechanisms, the query, key, and value projections play distinct roles in capturing relationships across the input sequence. If they are overly similar due to identical initialization, the model’s ability to attend to relevant and diverse information is compromised. One of the key strengths of self-attention lies in its capacity to model complex dependencies between different positions in the input. Therefore, when $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ share the same initialization, this capacity is diminished, leading to reduced performance. We further verified this observation using Swin-S by applying the DCT-based initialization to both $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$. The results confirmed the same trend, supporting our conclusion that single-matrix DCT initialization is generally more effective.
Removing DCT in Compressed Attention
This study discusses the necessity of DCT in the DCT-compressed attention. Specifically, removing DCT reduces the compression mechanism to a form of direct coefficient pruning in the spatial domain, without exploiting frequency-domain decorrelation. As Fig. 4b presents, when we removed the DCT in the DCT-compressed attention and retained other components, the accuracy of the revised Swin-T models on ImageNet-1K decreased across all compression ratios. These experiments show that DCT is essential because it transforms feature representations into the frequency domain, reducing redundancy and enhancing the model’s ability to capture essential information. As a result, the attention mechanism benefits from improved expressiveness and robustness, even under compression. The accuracy drop observed in Fig. 4b confirms that DCT plays a crucial role in preserving discriminative features, demonstrating its necessity in the proposed approach.
Backbone Param Object Detection Instance Segmentation
(M) $\text{AP}^{\text{Mask}}$ $\text{AP}^{\text{Mask}}_{0.5}$ $\text{AP}^{\text{Mask}}_{0.75}$ $\text{AP}^{\text{Mask}}$ $\text{AP}^{\text{Mask}}_{0.5}$ $\text{AP}^{\text{Mask}}_{0.75}$
Swin-T 86 50.3 69.1 54.3 43.7 66.6 47.3
DCFormer-T 86 43.5 62.6 47.4 38.0 59.3 40.7
@@
Table 5: COCO results.
4.4 COCO 2017 Benchmark
We further evaluated the proposed methods on COCO 2017 object detection and instance segmentation tasks Lin et al. (2014). We took Swin Transformers with ImageNet-1K pre-trained weights as the backbone and the Cascade Mask R-CNN Cai and Vasconcelos (2018) as the pipeline. Based on the superior performance observed in the ImageNet-1K experiments when initializing $\mathbf{W}_{K}$ with the DCT matrix, we adopted this initialization for the Swin-T and Swin-S backbones in our evaluation. As shown in Table 5, both models achieved slightly higher average precision (AP) scores compared to their vanilla counterparts across both tasks. We also evaluated the DCT-compressed attention variant with the 0.75 compression ratio. This modification allowed Swin-T to maintain comparable AP performance while reducing the parameter count by approximately 3 million. It also enabled Swin-S to achieve improved AP scores with around 5 million fewer parameters.
5 Conclusion
In this work, we introduced two new methods using DCT to decorrelate the input in the attention for vision transformers. First, we presented a novel DCT-based initialization approach to address the challenges associated with training attention weights. We demonstrated a significant improvement over traditional initialization, providing a robust foundation for the attention mechanism. Second, we formulated a DCT-based compression method that can effectively reduce the computational overhead by truncating high-frequency components. It resulted in smaller weight matrices that compromised considerable accuracy. We validated the efficacy of these breakthroughs, showcasing their potential to enhance the performance and efficiency of Transformer models in classification tasks. These contributions underscore the importance of thoughtful initialization strategies and compression techniques in advancing the capabilities of the Vision Transformer architectures.
Acknowledgments
H. Pan and U. Bagci were supported by NIH grants R01-HL171376 and U01-CA268808. E. Hamdan, X. Zhu, and A. E. Cetin were supported by NSF 2531376.